OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (3, 2) and (5, 12)

OpenStudy (anonymous):

This is how you figure out slope when given two points: (y2-y1)/(x2-x1)= rise/run (12-2)/(5-3) 10/2=5/1 or 5

OpenStudy (anonymous):

@Farmnat amazing thanks so much xxxxxxoooo